A 25 year old mother has a 5 year old daughter. In how many years will the mother be exactly twice as old as the daughter?
Found 2 solutions by stanbon, addingup:
Answer by stanbon(75887)

A 25 year old mother has a 5 year old daughter. In how many years will the mother be exactly twice as old as the daughter?
----
Let the # of years you need be "x"::
25+x = 2(5+x)
------------------
25+x = 10 + 2x
x = 15 years
---------------------
Cheers,
Stan H.
-------------

Answer by addingup(3677)

2(5+x) = 25+x Multiply on left
10+2x = 25+x Subtract 10 and x on both sides
x = 15 is your answer. In 15 years:
mother: 25+15= 40
daughter: 5+15 = 20 so mother is twice as old as the daughter.